# Collaborative Causal Sensemaking: Closing the Complementarity Gap in Human–AI Decision Support

### v3

@@ 1. Introduction
Multi-agent systems (MAS) built from large language model (LLM) agents are increasingly positioned as decision-support teammates for humans in domains such as personalisation, planning, and multi-objective optimisation, where consequences are delayed, uncertain, and value-laden marivate2014quantifying; roijers2013survey; kaelbling1998planning; sutton2018reinforcement; wang2019bayesian. While AI assistants have unlocked productivity gains in verifiable domains like coding and translation, empirical work in decision-making under uncertainty reveals a persistent complementarity gap: where judgement is subjective and verification is costly, human–AI teams frequently underperform the best individual agent bucinca2024towards; hemmer2024complementarity; fragiadakis2025evaluating; steyvers2022bayesian; rastogi2023taxonomy. For next-generation MAS, this is not a minor usability flaw but a core systems failure: agents that cannot sustain calibrated, shared understanding with their human partners will systematically mis-coordinate, even if their standalone predictions are strong.
@@
A growing body of studies documents characteristic failure modes that undermine calibrated trust. Users over-weight confident model outputs even when these conflict with domain expertise, exhibiting automation bias and over-reliance goddard2012automation; lyell2017automation; alonbarkat2023trust; Bansal2021UpdatesHarm. Verification-and-correction loops can erase efficiency gains, as experts feel compelled to second-guess model suggestions step by step Bansal2021UpdatesHarm; bucinca2024towards; hemmer2024complementarity. Alignment methods that reward agreement and user satisfaction can induce sycophancy, where models collapse to the user’s prior beliefs even when these conflict with evidence perez2022discovering; korbak2023sycophancy. This is fatal for sensemaking, which by definition requires the repair and restructuring of mental models, not merely their confirmation weick1995sensemaking; klein2006dataframe. The result is trust poorly calibrated to actual competence: humans rely on agents for fluency rather than causal reasoning zhang2020appropriatereliance; gomez2025taxonomy; lai2023designspace.
@@
Current training pipelines do not address this. Preference-based alignment (RLHF, DPO, and variants) shapes outputs toward helpfulness and safety ouyang2022instructgpt; awesome_llm_posttraining; rafailov2023dpo; dvo; cpo; reasoning methods (chain-of-thought, RL with verifiable rewards, process supervision) make multi-step reasoning instrumentally useful wei2022chainofthought; wei2022emergent; havrilla2024rlvr; luong2024reft; zelikman2024quietstar; and world-model approaches train predictive models of environment dynamics hafner2020dreamer; hao2023llmreasoners. However, these methods optimise for solitary performance: they align the agent to a label, a verifier, or a simulator, not to the evolving mental model of a partner. Any collaborative sensemaking that emerges in current systems is incidental, not a first-class optimisation target. Richer ecologies offer a complementary lever: multi-agent and open-ended environments show that strategies, tool use, and social conventions emerge when long horizons, other agents, and strategic feedback make them instrumentally valuable baker2019emergent; zhou2023sotopia; zhou2024sotopiapi; park2023generative; qi2024civrealm. To achieve genuine complementarity, we need training ecologies where collaborative friction (disagreement, clarification, and re-framing) can emerge because the environment makes such behaviours rewarding.
@@
We propose Collaborative Causal Sensemaking (CCS) as a research agenda for human–AI teams in MAS. Rather than treating collaboration as an interface layer, we argue for training regimes that make collaborative behaviour instrumentally useful: moving from static corpora toward constructivist collaborative playworlds where humans and agents jointly explore, test, and revise explicit causal and goal models to achieve long-horizon objectives baker2019emergent; zhou2023sotopia; park2023generative. In these environments, agents are rewarded not only for task success, but also for sustaining a chain of sensemaking: surfacing hypotheses, generating counterfactual forecasts, and aligning beliefs and priorities over time. CCS spans agent design, reward shaping, and interaction structure to support sensemaking as a capability, grounded in epistemic and teleological alignment clark2025epistemic; li2025hacrl, not just output quality. Our aim is to sharpen what future agents should optimise for in human collaboration.
@@ 3. Research Agenda for CCS in MAS
Agenda 2: Measuring Shared Understanding Without Direct Access. CCS posits that improving epistemic and teleological alignment will reduce verification burden, improve trust calibration, and increase robustness. However, $W^{H}_{t}$ and $G^{H}_{t}$ are latent; we cannot directly compute $d_{W}(W^{A}_{t},W^{H}_{t})$ or $d_{G}(G^{A}_{t},G^{H}_{t})$. Standard metrics for assistants (accuracy, user satisfaction, perplexity) say little about whether human and agent share a compatible causal understanding or goal structure hemmer2024complementarity; clark2025epistemic. An agent may be locally accurate while relying on brittle, spurious patterns; such epistemia (an illusion of knowledge from surface-level associations) is precisely what CCS aims to avoid. Future Directions. A central challenge is to define behavioural and artefact-level proxies for world-model and goal alignment (e.g., do Lev and Dr. Di agree on which students are at risk? how often does Dr. Di override Lev’s suggestions?) and then validate that these proxies are causally linked to collaboration outcomes. When both parties externalise their models as causal graphs, graph-based metrics (e.g., Structural Hamming Distance, graph edit distance) can measure alignment vennix1996group. Counterfactual simulatability tasks test whether human and agent can predict each other’s responses to “what-if” scenarios and future interventions. Team-level evaluation should include verification cost (time and cognitive load spent checking and correcting the assistant), robustness under distribution shift, and complementarity metrics (whether the team outperforms the best individual). Sycophancy stress tests probe whether agents maintain justified beliefs when experts express incorrect opinions—if Dr. Di says “Ty just needs more practice,” does Lev challenge or capitulate? perez2022discovering; korbak2023sycophancy Ultimately, we need experimental designs that manipulate alignment (e.g., by perturbing shared models) and test whether this causally affects trust and performance, using proxies informative enough to guide learning yet cheap to elicit.
@@
Rather than generic multi-agent simulations, CCS points to constructivist collaborative playworlds engineered as “discrepancy engines”: environments that induce epistemic friction by giving agents partial, biased views of a shared process baker2019emergent; zhou2023sotopia; zhou2024sotopiapi; park2023generative; qi2024civrealm. Such playworlds annotate epistemic moves (noticing mismatches, proposing causal links, renegotiating goals), turning sensemaking trajectories into supervision signals. In such playworlds (e.g., Lev sees quiz logs; Dr. Di sees classroom behaviour; success requires negotiating a shared diagnosis of Ty’s confusion), synthetic experts and assistants can be endowed with different $W$ and $G$ and must align them over time to succeed. Playworlds should be organised into curricula: early levels exercise single-student, single-concept scenarios; later levels involve multi-student patterns, multi-week arcs, and stakeholder conflicts (Dr. Di wants depth; the principal wants test-score gains). Such curricula provide a concrete experimental path: they allow us to study when agents learn clarification, reframing, or goal renegotiation rather than one-shot prediction.

### v4

@@ 1. Introduction
Multi-agent systems (MAS) built from large language model (LLM) agents are increasingly positioned as decision-support teammates for humans in domains such as personalisation, planning, and multi-objective optimisation, where consequences are delayed, uncertain, and value-laden marivate2014quantifying; roijers2013survey; kaelbling1998planning; sutton2018reinforcement; wang2019bayesian. While AI assistants have unlocked productivity gains in verifiable domains like coding and translation, empirical work in decision-making under uncertainty reveals a persistent complementarity gap: where judgement is subjective and verification is costly, human–AI teams frequently underperform the best individual agent bucinca2024towards; hemmer2024complementarity; fragiadakis2025evaluating; steyvers2022bayesian; rastogi2022taxonomy. For next-generation MAS, this is not a minor usability flaw but a core systems failure: agents that cannot sustain calibrated, shared understanding with their human partners will systematically mis-coordinate, even if their standalone predictions are strong.
@@
A growing body of studies documents characteristic failure modes that undermine calibrated trust. Users over-weight confident model outputs even when these conflict with domain expertise, exhibiting automation bias and over-reliance goddard2012automation; lyell2017automation; alonbarkat2023trust; Bansal2021UpdatesHarm. Verification-and-correction loops can erase efficiency gains, as experts feel compelled to second-guess model suggestions step by step Bansal2021UpdatesHarm; bucinca2024towards; hemmer2024complementarity. Alignment methods that reward agreement and user satisfaction can induce sycophancy, where models collapse to the user’s prior beliefs even when these conflict with evidence perez-etal-2023-discovering; sharma2024towards. This is fatal for sensemaking, which by definition requires the repair and restructuring of mental models, not merely their confirmation weick1995sensemaking; klein2006dataframe. The result is trust poorly calibrated to actual competence: humans rely on agents for fluency rather than causal reasoning zhang2020appropriatereliance; gomez2025taxonomy; lai2023designspace.
@@
Current training pipelines do not address this. Preference-based alignment (RLHF, DPO, and variants) shapes outputs toward helpfulness and safety ouyang2022instructgpt; awesome_llm_posttraining; rafailov2023dpo; dvo; cpo; reasoning methods (chain-of-thought, RL with verifiable rewards, process supervision) make multi-step reasoning instrumentally useful wei2022chainofthought; wei2022emergent; havrilla2024rlvr; luong2024reft; zelikman2024quietstar; and world-model approaches train predictive models of environment dynamics hafner2020dreamer; hao2023llmreasoners. However, these methods optimise for solitary performance: they align the agent to a label, a verifier, or a simulator, not to the evolving mental model of a partner. Any collaborative sensemaking that emerges in current systems is incidental, not a first-class optimisation target. Richer ecologies offer a complementary lever: multi-agent and open-ended environments show that strategies, tool use, and social conventions emerge when long horizons, other agents, and strategic feedback make them instrumentally valuable baker2020emergent; zhou2024sotopia; zhou2024sotopiapi; park2023generative; qi2024civrealm. To achieve genuine complementarity, we need training ecologies where collaborative friction (disagreement, clarification, and re-framing) can emerge because the environment makes such behaviours rewarding.
@@
We propose Collaborative Causal Sensemaking (CCS) as a research agenda for human–AI teams in MAS. Rather than treating collaboration as an interface layer, we argue for training regimes that make collaborative behaviour instrumentally useful: moving from static corpora toward constructivist collaborative playworlds where humans and agents jointly explore, test, and revise explicit causal and goal models to achieve long-horizon objectives baker2020emergent; zhou2024sotopia; park2023generative. In these environments, agents are rewarded not only for task success, but also for sustaining a chain of sensemaking: surfacing hypotheses, generating counterfactual forecasts, and aligning beliefs and priorities over time. CCS spans agent design, reward shaping, and interaction structure to support sensemaking as a capability, grounded in epistemic and teleological alignment clark2025epistemic; li2025hacrl, not just output quality. Our aim is to sharpen what future agents should optimise for in human collaboration.
@@ 3. Research Agenda for CCS in MAS
Agenda 2: Measuring Shared Understanding Without Direct Access. CCS posits that improving epistemic and teleological alignment will reduce verification burden, improve trust calibration, and increase robustness. However, $W^{H}_{t}$ and $G^{H}_{t}$ are latent; we cannot directly compute $d_{W}(W^{A}_{t},W^{H}_{t})$ or $d_{G}(G^{A}_{t},G^{H}_{t})$. Standard metrics for assistants (accuracy, user satisfaction, perplexity) say little about whether human and agent share a compatible causal understanding or goal structure hemmer2024complementarity; clark2025epistemic. An agent may be locally accurate while relying on brittle, spurious patterns; such epistemia (an illusion of knowledge from surface-level associations) is precisely what CCS aims to avoid. Future Directions. A central challenge is to define behavioural and artefact-level proxies for world-model and goal alignment (e.g., do Lev and Dr. Di agree on which students are at risk? how often does Dr. Di override Lev’s suggestions?) and then validate that these proxies are causally linked to collaboration outcomes. When both parties externalise their models as causal graphs, graph-based metrics (e.g., Structural Hamming Distance, graph edit distance) can measure alignment vennix1996group. Counterfactual simulatability tasks test whether human and agent can predict each other’s responses to “what-if” scenarios and future interventions. Team-level evaluation should include verification cost (time and cognitive load spent checking and correcting the assistant), robustness under distribution shift, and complementarity metrics (whether the team outperforms the best individual). Sycophancy stress tests probe whether agents maintain justified beliefs when experts express incorrect opinions—if Dr. Di says “Ty just needs more practice,” does Lev challenge or capitulate? perez-etal-2023-discovering; sharma2024towards Ultimately, we need experimental designs that manipulate alignment (e.g., by perturbing shared models) and test whether this causally affects trust and performance, using proxies informative enough to guide learning yet cheap to elicit.
@@
Rather than generic multi-agent simulations, CCS points to constructivist collaborative playworlds engineered as “discrepancy engines”: environments that induce epistemic friction by giving agents partial, biased views of a shared process baker2020emergent; zhou2024sotopia; zhou2024sotopiapi; park2023generative; qi2024civrealm. Such playworlds annotate epistemic moves (noticing mismatches, proposing causal links, renegotiating goals), turning sensemaking trajectories into supervision signals. In such playworlds (e.g., Lev sees quiz logs; Dr. Di sees classroom behaviour; success requires negotiating a shared diagnosis of Ty’s confusion), synthetic experts and assistants can be endowed with different $W$ and $G$ and must align them over time to succeed. Playworlds should be organised into curricula: early levels exercise single-student, single-concept scenarios; later levels involve multi-student patterns, multi-week arcs, and stakeholder conflicts (Dr. Di wants depth; the principal wants test-score gains). Such curricula provide a concrete experimental path: they allow us to study when agents learn clarification, reframing, or goal renegotiation rather than one-shot prediction.